# Exact fluctuation relation for open systems beyond the Jarzynski equality

### v1

@@ Abstract
We derive exact fluctuation equalities for open systems that recover free energy differences between two equilibrium endpoints connected by nonequilibrium processes with arbitrary dynamics and coupling. The exponential of the free energy difference is expressed in terms of ensemble averages of the Hamiltonian of mean force (HMF) shift and the chi-squared divergence between the initial and final marginal probability distribution of the open system. A trajectory counterpart of this relation follows from an asymptotic equilibration postulate, which treats relaxation to the final stationary canonical state as a boundary condition rather than as a consequence of constraints on the driven dynamics. In the frozen-coupling regime, the HMF shift reduces to the bare-system Hamiltonian shift, yielding a clear heat–work decomposition. The Jarzynski equality (JE) is recovered under the assumption of Hamiltonian dynamics for the combined system. We validate the theory on a dissipative, phase-space-compressing drive followed by an underdamped Langevin relaxation, where the assumptions underlying the JE break down, whereas our equality reproduces the exact free energy differences.
@@
The Jarzynski equality (JE) [1], $\langle e^{-\beta W}\rangle=e^{-\beta\Delta F}$, is a fundamental result in modern statistical mechanics. It establishes an exact connection between the exponential average of the nonequilibrium work $W$ performed during an irreversible process and the equilibrium free energy difference $\Delta F$. Along with related fluctuation theorems (FTs) [2, 3, 4], it has been confirmed in both classical and quantum systems and provides a powerful framework for extracting equilibrium information from nonequilibrium measurements [5, 6, 7, 8, 9, 10, 11, 12]. Nevertheless, its generality must be interpreted with care in different situations [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. The theoretical foundation of the original JE rests on strict dynamical reversibility [1]. In its original derivation, the composite system $\mathcal{S}+\mathcal{E}$, with bare system $\mathcal{S}$ and environment $\mathcal{E}$, evolves under deterministic dynamics that obey Liouville’s theorem, ensuring phase–space volume preservation. This preservation implies a fundamental dynamical constraint underlying the equality. A stochastic generalization soon followed, reformulating the dynamics through a Markov master equation or Langevin description [29]. In this generalization, reversibility takes a statistical form through detailed balance (DB). DB ensures that, for each fixed control parameter $\lambda$, the Gibbs–Boltzmann canonical distribution serves as the stationary state probability distribution, providing the probabilistic analogue of microscopic reversibility. The necessity of reversibility constraints was highlighted by Cohen and Mauzerall’s criticism [22], which questioned the formal validity of the JE and suggested its applicability might be limited to near-equilibrium weakly coupled regimes. Jarzynski’s reply [30] demonstrated that the JE is a mathematical identity following directly from Hamiltonian dynamics of the composite system $(\mathcal{S}+\mathcal{E})$. This derivation, later refined to incorporate strong coupling through the Hamiltonian of mean force (HMF) $H_{\beta}^{*}(X_{\mathcal{S}},\lambda)$, confirmed that while the free energy interpretation changes, the proof still relies on Liouville’s theorem and therefore on microscopic reversibility of the composite dynamics. Speck and Seifert [31] extended the equality to non-Markovian processes governed by the generalized Langevin equation. For such dynamics, the JE remains valid, provided the friction kernel and noise correlations obey the fluctuation–dissipation theorem (FDT). Compliance with FDT ensures that the Gibbs–Boltzmann distribution remains stationary for each frozen $\lambda$. While the JE demonstrated resilience against classical memory effects, a more fundamental challenge arose from an observation that the standard JE is necessarily violated in the presence of feedback — a phenomenon central to Maxwell’s Demon [32] and experimentally realized in information ratchets[33]. A formal resolution was achieved through the generalized Jarzynski equality (GJE) under feedback control, as formulated by Sagawa and Ueda [21, 14]. That extension incorporates the stochastic mutual information gained during measurement as a thermodynamic resource. However, this GJE still relies on a dynamical assumption, requiring the classical stochastic processes to satisfy local detailed balance (LDB) to ensure that the time-reversal property of path probabilities is well defined [34]. Regardless of the distinct approaches for the JE derivations and related generalizations, they all share an inherent dependency on underlying dynamical constraints such as microscopic reversibility, DB, FDT, or LDB [2, 35]. When these constraints are violated, the JE (or GJE) is no longer guaranteed to hold. Across biology, chemistry, physics, and engineering, many systems are driven under conditions that break these dynamical constraints, yet still relax to a genuine equilibrium once the driving is kept fixed. [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48].
The key question is whether it is possible to quantify the free energy difference between two equilibrium points connected by nonequilibrium processes, given arbitrary dynamics and system-environment coupling. In this Letter, we derive equalities for this most general scenario. As the first result, we derive an endpoint equality in terms of the driving and coupling protocols (Thm. 1), showing that its validity is independent of both the underlying dynamics and the strength of system–environment coupling linking the equilibrium endpoints. The second result (Thm. 2) establishes the trajectory counterpart based on the asymptotic postulate of thermal equilibration. For strongly coupled systems in the frozen–coupling regime — the physically most natural case — the equalities depend on the HMF in exactly the same way as on the bare–system Hamiltonian (Cor. 1). We then derive a heat–work decomposition (Cor. 2), which reveals the thermodynamic structure of the relations and provides a consistent interpretation of work and heat exchange. In the limit of Liouvillian dynamics and frozen coupling, our framework reduces to the standard JE. We validate the theory on a composite model, where the standard work estimator of JE fails under non-Liouvillian driving, while the endpoint and trajectory counterpart equalities reproduce the exact equilibrium free energy differences. We begin with a few necessary definitions (Def. 1-3) that later serve as the foundation for the endpoint and trajectory relations (Thm. 1 & 2) derived in this work.
Figure 1: (A) Schematic of the composite $\mathcal{S}+\mathcal{E}$ at a fixed bath temperature $T=\beta^{-1}/k_{B}$. The control $\lambda(t)$ (schematic purple spring) and the coupling $C(t)$ (schematic black spring) are driven independently and become constant at times $t_{\lambda}$ and $t_{C}$, respectively; the composite system then relaxes to equilibrium at $t_{\rm eq}$. The system free energy difference $\Delta F_{\mathcal{S}}^{*}(\beta)$ refers to the canonical endpoints $(\lambda(0),C(0))\!\to\!(\lambda(t_{\rm eq}),C(t_{\rm eq}))$ and is defined through the HMF partition functions. (B) An initial microstate $X_{0}\!\sim\!P_{0}(X,\beta)$ on $\Gamma(0)$ evolves under the kinematic map $X(t|X_{0})=\mathcal{T}_{t}(X_{0})$ through the ramp and the post–ramp relaxation, yielding the asymptotic density $P_{t_{\rm eq}}(X,\beta)$ on $\Gamma(t_{\rm eq})$. The frozen coupling case, $C(t)\equiv C(0)$, is the regime used in our validation.
Definition 1 (Microscopic trajectory of the composite system).
Here, we formulate the evolution of the composite system in an abstract manner, remaining agnostic to the details of its underlying dynamics. We consider a composite system $\mathcal{S}+\mathcal{E}$, consisting of a system of interest $\mathcal{S}$ and its environment $\mathcal{E}$. The total phase space of the composite at time $t$ is denoted by $\Gamma_{t}$, and a single microscopic state by $X\in\Gamma_{t}$. Each microstate contains the full phase-space coordinates of both parts, $X=(X_{\mathcal{S}},X_{\mathcal{E}})$, where $X_{\mathcal{S}}$ and $X_{\mathcal{E}}$ collect the positions and momenta of particles in $\mathcal{S}$ and $\mathcal{E}$, respectively. We introduce the trajectory map $\mathcal{T}_{t}:\Gamma_{0}\rightarrow\Gamma_{t}$, which assigns to each initial state $X_{0}\in\Gamma_{0}$, its evolved image $X(t|X_{0})=\mathcal{T}_{t}(X_{0})$ at time $t$. The map $\mathcal{T}_{t}$ is a purely kinematic construct and does not rely on any specific dynamical generator. It may represent Hamiltonian, stochastic, or arbitrary evolution, providing a unified description for all possible dynamics. Since $\mathcal{T}_{t}$ acts on the composite microstates $X_{0}\in\Gamma_{0}$, no assumption is made that the system and environment trajectories can be defined independently — a requirement intrinsic to any open system description. The microscopic state of the composite system along a trajectory is therefore written as $X(t|X_{0})=\bigl(X_{\mathcal{S}}(t|X_{0})=\mathcal{T}_{t}^{\mathcal{S}}(X_{0}),\,X_{\mathcal{E}}(t|X_{0})=\mathcal{T}_{t}^{\mathcal{E}}(X_{0})\bigr)$, where $\mathcal{T}_{t}^{\mathcal{S}}$ and $\mathcal{T}_{t}^{\mathcal{E}}$ denote the respective projections of the composite map $\mathcal{T}_{t}$ onto the system and environment coordinates. Next, we specify the Hamiltonian structure along the established trajectory.
Definition 2 (Total Hamiltonian along a trajectory).
The instantaneous total Hamiltonian of the composite system $\mathcal{S}+\mathcal{E}$ evaluated along the trajectory $\mathcal{T}_{t}(X_{0})$ is defined as
@@
$\quad+\mathcal{H}_{\mathcal{E}}\bigl(\mathcal{T}_{t}^{\mathcal{E}}(X_{0})\bigr)+\mathcal{V}_{\mathcal{S}\mathcal{E}}\bigl(\mathcal{T}_{t}^{\mathcal{S}}(X_{0}),\mathcal{T}_{t}^{\mathcal{E}}(X_{0}),C(t)\bigr).$ (1)
Here $\mathcal{H}_{\mathcal{S}}$ and $\mathcal{H}_{\mathcal{E}}$ are the bare Hamiltonians of the system and environment, respectively, and $\mathcal{V}_{\mathcal{S}\mathcal{E}}$ is their interaction potential. Two externally controlled parameters appear: $\lambda(t)$ represents the driving protocol acting on $\mathcal{S}$ (for example, a control parameter or mechanical coordinate) and $C(t)$ controls the coupling strength or interaction channel between $\mathcal{S}$ and $\mathcal{E}$. By allowing $\lambda(t)$ and $C(t)$ to vary independently, we retain the most general energetic structure of an open system subject to simultaneously applied driving and coupling controls. Each protocol may correspond to a sudden quench or to a continuous ramp in time. We denote by $t_{\lambda}$ and $t_{c}$ the times at which the corresponding protocols become constant, $\lambda(t)=\mathrm{const}\quad\text{for }t\geq t_{\lambda}$ and $C(t)=\mathrm{const}\quad\text{for }t\geq t_{c}$. The composite system subsequently relaxes to equilibrium at a later time $t_{\mathrm{eq}}$. In general, $0\leq t_{\lambda},\;t_{c}\leq t_{\mathrm{eq}}$ such that $t_{\lambda}\neq t_{c}$, so the completion of driving, coupling, and equilibration do not necessarily coincide. This distinction will later be crucial when we analyze the frozen–coupling limit $C(t)\equiv C(0)$. Now, we introduce the canonical ensembles for the composite system and the system of interest $\mathcal{S}$. These equilibrium measures serve as statistical reference points for the initial and final endpoints of the process.
Definition 3 (Canonical ensembles of the composite system and marginal distributions of the system).
The equilibrium probability distribution of the composite system $\mathcal{S}+\mathcal{E}$ at temperature $\beta^{-1}/k_{\mathrm{B}}$ is defined in canonical form as $P(X,\beta)={e^{-\beta\,\mathcal{H}_{\mathcal{S}+\mathcal{E}}(X,\lambda,C)}}/{\mathcal{Z}_{\mathcal{S}+\mathcal{E}}(\lambda,C,\beta)},$ where $\mathcal{H}_{\mathcal{S}+\mathcal{E}}$ is the total Hamiltonian given in Eq. (2) and $\mathcal{Z}_{\mathcal{S}+\mathcal{E}}$ denotes the partition function of the composite system, $\mathcal{Z}_{\mathcal{S}+\mathcal{E}}(\lambda,C,\beta)=\int_{\Gamma_{t}}\mathrm{d}X\,e^{-\beta\,\mathcal{H}_{\mathcal{S}+\mathcal{E}}(X,\lambda,C)}.$ Two specific equilibrium ensembles are relevant to the present discussion, i.e.
$P_{0}(X,\beta)$ $=\frac{e^{-\beta\,\mathcal{H}_{\mathcal{S}+\mathcal{E}}(X,\lambda(0),C(0))}}{\mathcal{Z}_{\mathcal{S}+\mathcal{E}}(\lambda(0),C(0),\beta)},$ (2)
$P_{t_{\mathrm{eq}}}(X,\beta)$ $=\frac{e^{-\beta\,\mathcal{H}_{\mathcal{S}+\mathcal{E}}(X,\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}))}}{\mathcal{Z}_{\mathcal{S}+\mathcal{E}}(\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}),\beta)}.$ (3)
The distributions $P_{0}$ and $P_{t_{\mathrm{eq}}}$ describe the initial and final equilibrium states of the composite system at the same bath temperature $\beta^{-1}/k_{\mathrm{B}}$. The equilibrium marginal distribution of the system $\mathcal{S}$ is obtained by integrating over the environmental coordinates $P^{\mathcal{S}}(X_{\mathcal{S}},\beta)=\int\mathrm{d}X_{\mathcal{E}}\,P(X_{\mathcal{S}},X_{\mathcal{E}},\beta)$. This probability can be expressed in canonical form through the HMF, defined as
@@
$\qquad-\frac{1}{\beta}\ln\int\mathrm{d}X_{\mathcal{E}}\,e^{-\beta\,[\mathcal{H}_{\mathcal{E}}(X_{\mathcal{E}})+\mathcal{V}_{\mathcal{S}\mathcal{E}}(X_{\mathcal{S}},X_{\mathcal{E}},C)]}.$ (4)
Equation (3) defines an effective Hamiltonian that incorporates the influence of the environment on $\mathcal{S}$ at fixed $(\lambda,C,\beta)$. The corresponding partition function of the system is $\mathcal{Z}_{\mathcal{S}}^{*}(\lambda,C,\beta)=\int\mathrm{d}X_{\mathcal{S}}\,e^{-\beta\,\mathcal{H}_{\beta}^{*}(X_{\mathcal{S}},\lambda,C)}$. Hence the equilibrium marginal distribution of the system reads $P^{\mathcal{S}}(X_{\mathcal{S}},\beta)={e^{-\beta\,\mathcal{H}_{\beta}^{*}(X_{\mathcal{S}},\lambda,C)}}/{\mathcal{Z}_{\mathcal{S}}^{*}(\lambda,C,\beta)}$. Applying them to the two equilibrium endpoints of the process yields
$P^{\mathcal{S}}_{0}(X_{\mathcal{S}},\beta)$ $=\frac{e^{-\beta\,\mathcal{H}_{\beta}^{*}(X_{\mathcal{S}},\lambda(0),C(0))}}{\mathcal{Z}_{\mathcal{S}}^{*}(\lambda(0),C(0),\beta)},$ (5)
$P^{\mathcal{S}}_{t_{\mathrm{eq}}}(X_{\mathcal{S}},\beta)$ $=\frac{e^{-\beta\,\mathcal{H}_{\beta}^{*}(X_{\mathcal{S}},\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}))}}{\mathcal{Z}_{\mathcal{S}}^{*}(\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}),\beta)}.$ (6)
Eqs. (5) and (6) define the initial and final equilibrium probability distributions of the system in terms of its HMF, corresponding to the protocol endpoints $(\lambda(0),C(0))$ and $(\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}))$, respectively. In Thm. 1, we show the exact connection between the system free energy difference at two endpoints and the canonical distributions and related HMFs at those endpoints.
Theorem 1 (Endpoint equalities for free energy differences).
For two equilibrium endpoints of the composite system $\mathcal{S}+\mathcal{E}$ prepared at the same inverse temperature $\beta$, we have (see Appendix for the proof)
@@
The free energy difference is
$\Delta F^{*}_{\mathcal{S}}(\beta)=F^{*}_{\mathcal{S}}(\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}),\beta)-F^{*}_{\mathcal{S}}(\lambda(0),C(0),\beta),$ (9)
where $F^{*}_{\mathcal{S}}(\lambda,C,\beta)=\beta^{-1}\ln\mathcal{Z}_{\mathcal{S}}^{*}(\lambda,C,\beta)$ and the HMF shift is
$\Delta\mathcal{H}^{*}_{\beta}(X_{\mathcal{S}})=\mathcal{H}^{*}_{\beta}(X_{\mathcal{S}},\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}))$
$\qquad-\mathcal{H}^{*}_{\beta}(X_{\mathcal{S}},\lambda(0),C(0)).$ (10)
Averages are taken over the final equilibrium ensemble,
@@
and the chi–squared divergence between the endpoint marginals is
$1+\chi^{2}\bigl(P^{\mathcal{S}}_{t_{\mathrm{eq}}}\parallel P^{\mathcal{S}}_{0}\bigr)=\int\mathrm{d}X_{\mathcal{S}}\,\frac{\bigl(P^{\mathcal{S}}_{t_{\mathrm{eq}}}(X_{\mathcal{S}},\beta)\bigr)^{2}}{P^{\mathcal{S}}_{0}(X_{\mathcal{S}},\beta)}.$ (12)
The result established here coincides with the endpoint equality derived in  [49], where the protocol is implicitly present. As can be seen from the derivation of Eqs. (7) and (8), no restriction is imposed on either the coupling or the form of the dynamics. The reasoning relies only on the existence of well-defined equilibrium endpoints, which guarantees that the corresponding marginal distributions and free energy difference are statistically and thermodynamically meaningful, respectively. Before introducing the trajectory counterparts, it is essential to clarify the assumption underlying their derivation. We postulate asymptotic equilibration of the full probability distribution along the trajectory. This assumption expresses the natural requirement that the Helmholtz free energy difference $\Delta F$ be well defined from the trajectory perspective of any nonequilibrium process connecting two equilibrium states, without imposing any particular dynamical constraint during the driving stage. It contrasts sharply with the standard derivations of the JE and its extensions, where the equality holds only if the underlying dynamics, during the application of the driving protocol, evolve toward a canonical stationary state once the control parameters are held fixed. In other words, the stationary canonical distribution at equilibrium is a consequence of constraints on the driven dynamics. By treating asymptotic convergence to $P_{t_{\mathrm{eq}}}$ as a natural boundary condition of thermal contact — rather than as a consequence of these dynamical constraints — the definition of equilibrium becomes decoupled from the specific generator of dynamics during the driving stage. This decoupling extends the application of our derivations to a broader class of systems, including those governed by arbitrary dynamics that are incompatible with the underlying assumptions of JE. We will later discuss the subtleties surrounding the imposition of that boundary condition in the validation section.
@@
Using the trajectory map $\mathcal{T}_{t}:\Gamma_{0}\rightarrow\Gamma_{t}$ defined in Def. 1, the time evolution of the probability density of the composite system can be defined as
$P_{t}(X,\beta)=\int_{\Gamma_{0}}\mathrm{d}X_{0}\,\delta\bigl(X-\mathcal{T}_{t}(X_{0})\bigr)\,P_{0}(X_{0},\beta),$ (13)
where $P_{0}(X_{0},\beta)$ is the initial canonical distribution given in Eq. (2). Integrating over the environmental degrees of freedom yields the time evolution of the system’s marginal distribution (see Appendix for derivation)
$P_{t}^{\mathcal{S}}(X_{\mathcal{S}},\beta)=\int_{\Gamma_{0}}\mathrm{d}X_{0}\,P_{0}(X_{0},\beta)\,\delta\bigl(X_{\mathcal{S}}-\mathcal{T}_{t}^{\mathcal{S}}(X_{0})\bigr).$ (14)
We postulate the asymptotic equilibration of the full distribution that implies convergence of the time evolution of the system’s marginal distribution to its canonical endpoint form
$\lim_{t\rightarrow t_{\mathrm{eq}}}P_{t}^{\mathcal{S}}(X_{\mathcal{S}},\beta)=P^{\mathcal{S}}_{t_{\mathrm{eq}}}(X_{\mathcal{S}},\beta).$ (15)
Under that condition, the ensemble averages in Thm. 1 admit the exact trajectory representation (see Appendix for proof), with averages over the initial canonical ensemble, $\langle\bullet\rangle_{X_{0}}=\int\mathrm{d}X_{0}\,\bullet\,P_{0}(X_{0},\beta)$. Consequently, the endpoint equalities (7) and (8) take the trajectory form
$e^{-\beta\Delta F^{*}_{\mathcal{S}}(\beta)}=\frac{\bigl\langle e^{-\beta\Delta\mathcal{H}^{*}_{\beta}\bigl(\mathcal{T}_{t_{\mathrm{eq}}}^{\mathcal{S}}(X_{0})\bigr)}\bigr\rangle_{X_{0}}}{1+\chi^{2}\bigl(P^{\mathcal{S}}_{t_{\mathrm{eq}}}\parallel P^{\mathcal{S}}_{0}\bigr)},$ (16)

### v2

@@ Abstract
We derive exact fluctuation equalities for open systems that recover free energy differences between two equilibrium endpoints connected by nonequilibrium processes with arbitrary dynamics and coupling. The exponential of the free energy difference is expressed in terms of ensemble averages of the Hamiltonian of mean force (HMF) shift and the chi-squared divergence between the initial and final marginal probability distribution of the open system. A trajectory counterpart of this relation follows from an asymptotic equilibration postulate, which treats relaxation to the final stationary canonical state as a boundary condition rather than as a consequence of constraints on the driven dynamics. In the frozen-coupling regime, the HMF shift reduces to the bare-system Hamiltonian shift, yielding a clear heat-work-feedback decomposition. The Jarzynski equality (JE) is recovered under the assumption of Hamiltonian dynamics for the combined system. We validate the theory on a dissipative, phase-space-compressing drive followed by an underdamped Langevin relaxation, where the assumptions underlying the JE break down, whereas our equality reproduces the exact free energy differences.
@@
The Jarzynski equality (JE) [jarzynski1997nonequilibrium], $\langle e^{-\beta W}\rangle=e^{-\beta\Delta F}$, is a fundamental result in modern statistical mechanics. It establishes an exact connection between the exponential average of the nonequilibrium work $W$ performed during an irreversible process and the equilibrium free energy difference $\Delta F$. Along with related fluctuation theorems (FTs) [esposito2009nonequilibrium, crooks1998nonequilibrium, crooks1999entropy], it has been confirmed in both classical and quantum systems and provides a powerful framework for extracting equilibrium information from nonequilibrium measurements [liphardt2002equilibrium, ohzeki2010quantum, xiong2018experimental, liu2023experimental, harris2007experimental, douarche2005experimental, saira2012test, wimsatt2021harnessing]. Nevertheless, its generality must be interpreted with care in different situations [gore2003bias, toyabe2010experimental, deng2017deformed, hernandez2023experimental, hahn2023quantum, presse2006ordering, chen2008crooks, crooks2009comment, sagawa2010generalized, cohen2004note, cohen2005jarzynski, argun2016experimental, vilar2008failure, palmieri2007jarzynski, monge2018experimental, sone2020quantum]. The theoretical foundation of the original JE rests on strict dynamical reversibility [jarzynski1997nonequilibrium]. In its original derivation, the composite system $\mathcal{S}+\mathcal{E}$, with bare system $\mathcal{S}$ and environment $\mathcal{E}$, evolves under deterministic dynamics that obey Liouville’s theorem, ensuring phase-space volume preservation. This preservation implies a fundamental dynamical constraint underlying the equality. A stochastic generalization soon followed, reformulating the dynamics through a Markov master equation or Langevin description [jarzynski1997equilibrium]. In this generalization, reversibility takes a statistical form through detailed balance (DB). DB ensures that, for each fixed control parameter $\lambda$, the Gibbs-Boltzmann canonical distribution serves as the stationary state probability distribution, providing the probabilistic analogue of microscopic reversibility. The necessity of reversibility constraints was highlighted by Cohen and Mauzerall’s criticism [cohen2004note], which questioned the formal validity of the JE and suggested its applicability might be limited to near-equilibrium weakly coupled regimes. Jarzynski’s reply [jarzynski2004nonequilibrium] demonstrated that the JE is a mathematical identity following directly from Hamiltonian dynamics of the composite system $(\mathcal{S}+\mathcal{E})$. This derivation, later refined to incorporate strong coupling through the Hamiltonian of mean force (HMF) $H_{\beta}^{*}(X_{\mathcal{S}},\lambda)$, confirmed that while the free energy interpretation changes, the proof still relies on Liouville’s theorem and therefore on microscopic reversibility of the composite dynamics. Speck and Seifert [speck2007jarzynski] extended the equality to non-Markovian processes governed by the generalized Langevin equation. For such dynamics, the JE remains valid, provided the friction kernel and noise correlations obey the fluctuation-dissipation theorem (FDT). Compliance with FDT ensures that the Gibbs-Boltzmann distribution remains stationary for each frozen $\lambda$. While the JE demonstrated resilience against classical memory effects, a more fundamental challenge arose from an observation that the standard JE is necessarily violated in the presence of feedback — a phenomenon central to Maxwell’s Demon [maruyama2009colloquium] and experimentally realized in information ratchets [serreli2007molecular]. A formal resolution was achieved through the generalized Jarzynski equality (GJE) under feedback control, as formulated by Sagawa and Ueda [sagawa2010generalized, toyabe2010experimental]. That extension incorporates the stochastic mutual information gained during measurement as a thermodynamic resource. However, this GJE still relies on a dynamical assumption, requiring the classical stochastic processes to satisfy local detailed balance (LDB) to ensure that the time-reversal property of path probabilities is well defined [maes2021local]. Regardless of the distinct approaches for the JE derivations and related generalizations, they all share an inherent dependency on underlying dynamical constraints such as microscopic reversibility, DB, FDT, or LDB [esposito2009nonequilibrium, scholl2006proof]. When these constraints are violated, the JE (or GJE) is no longer guaranteed to hold. Across biology, chemistry, physics, and engineering, many systems are driven under conditions that break these dynamical constraints, yet still relax to a genuine equilibrium once the driving is kept fixed. [gaspard2020active, martinez2019inferring, gnesotto2018broken, goswami2019work, godec2022challenges, krishnamurthy2016micrometre, reimann2002brownian, bechinger2016active, caprini2019activity, vicsek2012collective, cates2015motility, fodor2016far, argun2016non]. The key question is whether it is possible to quantify the free energy difference between two equilibrium points connected by nonequilibrium processes, given arbitrary dynamics and system-environment coupling. A detailed treatment is given in a companion article [PRE], here we only recall the main setting needed for the present Letter. We consider a composite system $\mathcal{S}+\mathcal{E}$ with microscopic states $X=(X_{\mathcal{S}},X_{\mathcal{E}})\in\Gamma_{t}$. Its evolution is described by a trajectory map $\mathcal{T}_{t}:\Gamma_{0}\to\Gamma_{t}$,
$X(t|X_{0})=\mathcal{T}_{t}(X_{0}),$ (1)
with projections $\mathcal{T}_{t}^{\mathcal{S}}$, $\mathcal{T}_{t}^{\mathcal{E}}$ onto system and environment coordinates. The map $\mathcal{T}_{t}$ is purely kinematic and may represent Hamiltonian, stochastic, or more general dynamics; no specific generator is imposed (see [PRE] for details). Along a trajectory the total Hamiltonian is
@@
$\quad+\mathcal{H}_{\mathcal{E}}\bigl(\mathcal{T}_{t}^{\mathcal{E}}(X_{0})\bigr)+\mathcal{V}_{\mathcal{S}\mathcal{E}}\bigl(\mathcal{T}_{t}^{\mathcal{S}}(X_{0}),\mathcal{T}_{t}^{\mathcal{E}}(X_{0}),C(t)\bigr),$ (2)
where $\mathcal{H}_{\mathcal{S}}$, $\mathcal{H}_{\mathcal{E}}$, and $\mathcal{V}_{\mathcal{S}\mathcal{E}}$ are the bare system, environment, and interaction contributions. The protocol $\lambda(t)$ drives $\mathcal{S}$, while $C(t)$ controls its coupling to $\mathcal{E}$. We denote by $t_{\lambda}$ and $t_{c}$ the times after which $\lambda(t)$ and $C(t)$ are fixed, and by $t_{\mathrm{eq}}$ the later time at which the composite has relaxed to equilibrium at $(\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}))$. At initial and final fixed $(\lambda,C,\beta)$ the composite is in the canonical state
$P(X;\lambda,C,\beta)=\frac{e^{-\beta\,\mathcal{H}_{\mathcal{S}+\mathcal{E}}(X,\lambda,C)}}{\mathcal{Z}_{\mathcal{S}+\mathcal{E}}(\lambda,C,\beta)},$ (3)
with partition function
$\mathcal{Z}_{\mathcal{S}+\mathcal{E}}(\lambda,C,\beta)=\int_{\Gamma_{t}}\mathrm{d}X\,e^{-\beta\,\mathcal{H}_{\mathcal{S}+\mathcal{E}}(X,\lambda,C)}.$ (4)
The two equilibrium endpoints are
$P_{0}(X,\beta)$ $=P\bigl(X;\lambda(0),C(0),\beta\bigr),$ (5)
$P_{t_{\mathrm{eq}}}(X,\beta)$ $=P\bigl(X;\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}),\beta\bigr).$ (6)
The equilibrium marginal of the system is
$P^{\mathcal{S}}(X_{\mathcal{S}};\lambda,C,\beta)=\int\mathrm{d}X_{\mathcal{E}}\,P(X_{\mathcal{S}},X_{\mathcal{E}};\lambda,C,\beta).$ (7)
It admits a canonical representation through the Hamiltonian of mean force (HMF)
@@
$-\frac{1}{\beta}\ln\int\mathrm{d}X_{\mathcal{E}}\,\frac{1}{\mathcal{Z}_{\mathcal{E}}}e^{-\beta\,[\mathcal{H}_{\mathcal{E}}(X_{\mathcal{E}})+\mathcal{V}_{\mathcal{S}\mathcal{E}}(X_{\mathcal{S}},X_{\mathcal{E}},C)]},$ (8)
with partition function
$\mathcal{Z}_{\mathcal{S}}^{*}(\lambda,C,\beta)=\int\mathrm{d}X_{\mathcal{S}}\,e^{-\beta\,\mathcal{H}_{\beta}^{*}(X_{\mathcal{S}},\lambda,C)}$ (9)
and marginal
$P^{\mathcal{S}}(X_{\mathcal{S}};\lambda,C,\beta)=\frac{e^{-\beta\,\mathcal{H}_{\beta}^{*}(X_{\mathcal{S}},\lambda,C)}}{\mathcal{Z}_{\mathcal{S}}^{*}(\lambda,C,\beta)}.$ (10)
Applied to the endpoints, this yields
$P^{\mathcal{S}}_{0}(X_{\mathcal{S}},\beta)$ $=P^{\mathcal{S}}(X_{\mathcal{S}};\lambda(0),C(0),\beta),$ (11)
$P^{\mathcal{S}}_{t_{\mathrm{eq}}}(X_{\mathcal{S}},\beta)$ $=P^{\mathcal{S}}(X_{\mathcal{S}};\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}),\beta).$ (12)
The HMF free energy
$F^{*}_{\mathcal{S}}(\lambda,C,\beta)=-\beta^{-1}\ln\mathcal{Z}_{\mathcal{S}}^{*}(\lambda,C,\beta)$ (13)
defines the free energy difference
$\Delta F^{*}_{\mathcal{S}}(\beta)=F^{*}_{\mathcal{S}}(\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}),\beta)-F^{*}_{\mathcal{S}}(\lambda(0),C(0),\beta).$ (14)
Theorem 1 (Endpoint equalities for HMF free energy).
For two equilibrium endpoints at the same inverse temperature $\beta$,
@@
Here
@@
the HMF shift is $\Delta\mathcal{H}^{*}_{\beta}(X_{\mathcal{S}})=\mathcal{H}^{*}_{\beta}(X_{\mathcal{S}},\lambda(t_{\mathrm{eq}}),C(t_{\mathrm{eq}}))-\mathcal{H}^{*}_{\beta}(X_{\mathcal{S}},\lambda(0),C(0))$, and the chi-squared divergence between endpoint marginals is
$\chi^{2}\bigl(P^{\mathcal{S}}_{t_{\mathrm{eq}}}\parallel P^{\mathcal{S}}_{0}\bigr)=-1+\int\mathrm{d}X_{\mathcal{S}}\,\frac{\bigl(P^{\mathcal{S}}_{t_{\mathrm{eq}}}(X_{\mathcal{S}},\beta)\bigr)^{2}}{P^{\mathcal{S}}_{0}(X_{\mathcal{S}},\beta)}.$ (18)
The result established here coincides with the endpoint equality derived in [rahbar2025thermodynamic], where the protocol is implicitly present. The trajectory map $\mathcal{T}_{t}$ pushes the initial distribution $P_{0}$ forward to a family $P_{t}$, and its system projection defines the time-dependent marginal $P_{t}^{\mathcal{S}}$. We postulate asymptotic equilibration,
$\lim_{t\rightarrow t_{\mathrm{eq}}}P_{t}^{\mathcal{S}}(X_{\mathcal{S}},\beta)=P^{\mathcal{S}}_{t_{\mathrm{eq}}}(X_{\mathcal{S}},\beta),$ (19)
which guarantees that $\Delta F^{*}_{\mathcal{S}}$ is well defined for any nonequilibrium protocol connecting the two equilibrium endpoints, without constraining the driven dynamics to be microscopically reversible. The explicit pushforward expressions and a detailed discussion of this boundary condition are given in [PRE]. Under asymptotic equilibration, the endpoint equalities admit exact trajectory forms. Introducing the average over the initial canonical ensemble,
$\langle\bullet\rangle_{X_{0}}=\int\mathrm{d}X_{0}\,\bullet\,P_{0}(X_{0},\beta),$ (20)
we can obtain the following theorem.
@@
The HMF free energy difference satisfies
$e^{-\beta\Delta F^{*}_{\mathcal{S}}(\beta)}=\frac{\Big\langle e^{-\beta\Delta\mathcal{H}^{*}_{\beta}\bigl(\mathcal{T}_{t_{\mathrm{eq}}}^{\mathcal{S}}(X_{0})\bigr)}\Big\rangle_{X_{0}}}{1+\chi^{2}\bigl(P^{\mathcal{S}}_{t_{\mathrm{eq}}}\parallel P^{\mathcal{S}}_{0}\bigr)},$ (21)